mathslover (mathslover):

find the next number : 6,2,30,10,56,26,132, ? (only answers please)

OpenStudy (lgbasallote):

96

mathslover (mathslover):

oops wrong

OpenStudy (lgbasallote):

ahh so it really is -68 huh

mathslover (mathslover):

no

OpenStudy (lgbasallote):

lol

mathslover (mathslover):

IT IS THE WORST RIDDLE SORRY

mathslover (mathslover):

@ParthKohli

OpenStudy (shubhamsrg):

182 ?

hartnn (hartnn):

any hint ?

mathslover (mathslover):

yes... hint .. is "alternate numbers"

hartnn (hartnn):

that was the 1st thing i did,broke it into 2 series......still stuck.

mathslover (mathslover):

squares + 1 .. something like that in 1 series

Parth (parthkohli):

42

hartnn (hartnn):

so it would be 50 right ?49+1

hartnn (hartnn):

isn't it 50 ? @mathslover

hartnn (hartnn):

and 132=11^2+11,56=7^2+7,30=5^2+5.so 1st term should be 12=3^2+3....

hartnn (hartnn):

then u'll have 2^2+2=6

OpenStudy (shubhamsrg):

yep..or in other words,,its equilvalent to 11*12 , 5*6 , 7*8 ...

OpenStudy (anonymous):

@mathslover is 92 the answer

mathslover (mathslover):

2^2 +2 , 1^2+1 , 5^2 +5 , 3^2+1, 11^2+11 , 5^2 + 1 hence the answer is 26

mathslover (mathslover):

I think this was too much complicated and 'bad' I will post better one soon

Parth (parthkohli):

Yeah, it really was much too complicated. =/

hartnn (hartnn):

but 26 was already there ....so next would be 7^2+1 = 50 ,isn't it?